OpenStudy (anonymous):

A sound at 50 dB exerts a pressure that is _______ times greater than a sound at 30 dB. A.10 B.100 C.1000 D.10,000 2. The colors of visible light each correspond to a different frequency of electromagnetic waves. True False

OpenStudy (anonymous):

for number 2 violet: 668–789 THz indigo: 668–714 THz blue: 606–668 THz green: 526–606 THz yellow: 508–526 THz orange: 484–508 THz red: 400–484 THz

OpenStudy (anonymous):

I think db are in log I wanna say 100 but dont take me to serious in this one.

OpenStudy (kropot72):

\[20=20\times \log_{10} \frac{P _{1}}{P _{2}}\] \[\log_{10} \frac{P _{1}}{P _{2}}=1\] \[\frac{P _{1}}{P _{2}}=10\]

OpenStudy (anonymous):

thank you but i dont understand the answer to the first one

OpenStudy (anonymous):

Im Sorry i mean the second one .,!

OpenStudy (anonymous):

30db= 0.000632456 50db=0.006324555 10 is the correct answer. http://www.sengpielaudio.com/calculator-soundlevel.htm

OpenStudy (kropot72):

The dB is a logarithmic ratio. If the ratio of two sound pressure levels is 1000 to 1 then the ratio in dB is\[20\times \log_{10}\frac{1000}{1}=20\times 3=60dB\]

OpenStudy (kropot72):

Brent0423 the dB is a ratio not a value. Please explain your numbers

OpenStudy (anonymous):

those are the numbers for the pressure

OpenStudy (kropot72):

The numbers must be relative to a reference sound pressure. This reference level should be stated.